# captionUnknown document class (or package)

### v1

@@ IV-A Experimental Setup
MoE Model. Our experiments employ two MoE architectures: (1) Mixtral-8x7B [2] (32 layers with 2 active experts out of 8 per layer), whose complete parameter set exceeds single-GPU memory capacity, and (2) Deepseek-V2-Lite [15] (26 layers selecting 8 active experts from 64 per layer, totaling 1,664 possible expert combinations), which presents greater routing complexity while being more memory-efficient. To simulate realistic edge deployment scenarios, we artificially constrain GPU memory allocation to 70% of total capacity for Mixtral-8x7B and 30% for Deepseek-V2-Lite, reflecting their respective memory requirements and operational constraints.
Dataset. Our experimental evaluation incorporated four widely adopted benchmark datasets: BIG-bench [11], MMLU-Pro [16], WikiText [17], and Tako [18].
@@
  • Uniform. Experts are uniformly distributed across all available devices. For instance, in our experimental setup with 4 GPUs, each device hosts 2 random experts per layer for the Mixtral model and 16 experts per layer for Deepseek-V2-Lite. This placement method aligns with the expert parallelism implementation in the Megatron-LM [9] project.
@@
  • SmartMoE [19]. While SmartMoE’s core focus is training optimization, our framework specifically implements SmartMoE’s placement module, which strategically distributes experts across GPUs based on real-time workload. The placement algorithm balances workload distribution to maintain uniform expert allocation across devices.
  • Expert Parallelism Load Balancer (EPLB) [10]. EPLB is a redundant expert strategy in DeepSeek-V3 [3] that duplicates high-load experts and heuristically distributes them to balance GPU workloads. Since the open-source implementation only supports homogeneous systems, we re-implement the algorithm within our framework to accommodate heterogeneous system configurations.
@@ IV-B System Performance
Effectiveness of Migration. We evaluated migration effectiveness by comparing adaptive and static systems using Deepseek-V2-Lite (Fig. 7), from 200 MultiData to 200 BIG-bench requests per server. Both systems showed identical initial performance, but the migration-enabled approach (“w” in Figure 7(a)) achieved significantly higher local computation after the first migration. We now explain the three migrations shown in Figure 7(a). The first one is triggered because the system detects a divergence between the accumulated inference data and the dataset used for the initial placement, activating the condition in Eq.(4). The second migration follows an inference data shift (see Data Change in Figure 7(b)). However, since the cost $C(\cdot)$ in Eq.(4) calculated at this time is based on the accumulated inference data mixing the old and new data, a third migration occurs later as more new data accumulates. The migration version achieved a 10% reduction in average latency (7.48 to 6.73) of all requests compared to the non-migration.
@@ V Related Work
Several edge serving works [20, 21, 22] have explored efficient strategies to achieve low-latency application services, yet they fail to account for the unique architectural characteristics and parallelization patterns inherent in MoE models. Recent efforts to optimize Mixture-of-Experts (MoE) models span both training and inference phases, targeting challenges such as expert placement, memory efficiency, and communication overhead. For training scenarios, FasterMoE [23] optimizes training by trading off hidden-state transfers against expert migrations, while SmartMoE [19], FlexMoE [24] and Prophet [25] develop GPU load balancing strategies. Lazarus [26] further considers system-level scalability. However, these training-oriented approaches fail to address the heterogeneous communication and computational capabilities inherent in multi-server inference environments. For inference on memory-constrained devices, standalone systems have explored multi-level memory hierarchies and expert prefetching strategies. Methods include storing excess experts in CPU memory (as in Pre-gated MoE [27], which introduces pre-gating for parallel loading at potential accuracy cost), predicting expert activations to prefetch parameters (Lina [28], MoE-Infinity [29], EdgeMoE [30], AdapMoE [31], SwapMoE [32], SIDA [33]), and computing experts directly on CPU (Fiddler [34]). While effective for individual resource-constrained devices, these solutions neglect cross-server resource coordination. In contrast, DanceMoE systematically coordinates distributed resources through optimized collaboration.
More recently, inference-time expert placement has been explored in systems like Moetuner [8] and DeepSeek-V3’s EPLB [10]. These methods use historical routing or expert redundancy to balance loads, but are generally designed for cluster-scale datacenter deployment, without accounting for edge-specific constraints like heterogeneous hardware or limited bandwidth. Complementary to placement strategies, communication-aware MoE systems, such as PipeMoE [35], ScheMoE [36], and TUTEL [37], aim to improve system efficiency through pipelining, hybrid parallelism, or fine-grained communication scheduling. While primarily designed for training, their techniques offer complementary benefits for distributed inference. In contrast to these prior efforts, DanceMoE targets collaborative inference across memory-limited edge servers. It introduces activation-aware placement and lightweight migration strategies tailored to heterogeneous devices and variable workloads—an underexplored yet increasingly important setting for scalable MoE deployment.

### v2

@@ IV-A Experimental Setup
MoE Model. Our experiments employ two MoE architectures: (1) Mixtral-8x7B [jiang2024mixtral] (32 layers with 2 active experts out of 8 per layer), whose complete parameter set exceeds single-GPU memory capacity, and (2) Deepseek-V2-Lite [liu2024deepseekv2] (26 layers selecting 8 active experts from 64 per layer, totaling 1,664 possible expert combinations), which presents greater routing complexity while being more memory-efficient. To simulate realistic edge deployment scenarios, we artificially constrain GPU memory allocation to 70% of total capacity for Mixtral-8x7B and 30% for Deepseek-V2-Lite, reflecting their respective memory requirements and operational constraints.
Dataset. Our experimental evaluation incorporated four widely adopted benchmark datasets: BIG-bench [srivastava2023beyond], MMLU-Pro [wang2024mmlu], WikiText [merity2016pointer], and Tako [li2023taco].
@@
  • Uniform. Experts are uniformly distributed across all available devices. For instance, in our experimental setup with 4 GPUs, each device hosts 2 random experts per layer for the Mixtral model and 16 experts per layer for Deepseek-V2-Lite. This placement method aligns with the expert parallelism implementation in the Megatron-LM [shoeybi2019megatron] project.
@@
  • SmartMoE [zhai2023smartmoe]. While SmartMoE’s core focus is training optimization, our framework specifically implements SmartMoE’s placement module, which strategically distributes experts across GPUs based on real-time workload. The placement algorithm balances workload distribution to maintain uniform expert allocation across devices.
  • Expert Parallelism Load Balancer (EPLB) [deepseek2025eplb]. EPLB is a redundant expert strategy in DeepSeek-V3 [liu2024deepseek] that duplicates high-load experts and heuristically distributes them to balance GPU workloads. Since the open-source implementation only supports homogeneous systems, we re-implement the algorithm within our framework to accommodate heterogeneous system configurations.
@@ IV-B System Performance
Effectiveness of Migration. We evaluated migration effectiveness by comparing adaptive and static systems using Deepseek-V2-Lite (Fig. 7), from 200 MultiData to 200 BIG-bench requests per server. Both systems showed identical initial performance, but the migration-enabled approach (“w” in Figure 7(a)) achieved significantly higher local computation after the first migration. We now explain the three migrations shown in Figure 7(a). The first one is triggered because the system detects a divergence between the accumulated inference data and the dataset used for the initial placement, activating the condition in Eq.(4). The second migration follows an inference data shift (see Data Change in Figure 7(b)). However, since the cost $C(\cdot)$ in Eq.(4) calculated at this time is based on the accumulated inference data mixing the old and new data, a third migration occurs later as more new data accumulates.
The migration version achieved a 10% reduction in average latency (7.48 to 6.73) of all requests compared to the non-migration.
@@ V Related Work
Several edge serving works [li2022joint, li2023task, li2023tapfinger] have explored efficient strategies to achieve low-latency application services, yet they fail to account for the unique architectural characteristics and parallelization patterns inherent in MoE models. Recent efforts to optimize Mixture-of-Experts (MoE) models span both training and inference phases, targeting challenges such as expert placement, memory efficiency, and communication overhead. For training scenarios, FasterMoE [he2022fastermoe] optimizes training by trading off hidden-state transfers against expert migrations, while SmartMoE [zhai2023smartmoe], FlexMoE [nie2023flexmoe] and Prophet [wang2023prophet] develop GPU load balancing strategies. Lazarus [wu2024lazarus] further considers system-level scalability. However, these training-oriented approaches fail to address the heterogeneous communication and computational capabilities inherent in multi-server inference environments. For inference on memory-constrained devices, standalone systems have explored multi-level memory hierarchies and expert prefetching strategies. Methods include storing excess experts in CPU memory (as in Pre-gated MoE [hwang2024pre], which introduces pre-gating for parallel loading at potential accuracy cost), predicting expert activations to prefetch parameters (Lina [li2023accelerating], MoE-Infinity [xue2024moe], EdgeMoE [yi2023edgemoe], AdapMoE [zhong2024adapmoe], SwapMoE [kong2024swapmoe], SIDA [du2024sida]), and computing experts directly on CPU (Fiddler [kamahori2024fiddler]). While effective for individual resource-constrained devices, these solutions neglect cross-server resource coordination. In contrast, DanceMoE systematically coordinates distributed resources through optimized collaboration.
More recently, inference-time expert placement has been explored in systems like Moetuner [go2025moetuner] and DeepSeek-V3’s EPLB [deepseek2025eplb]. These methods use historical routing or expert redundancy to balance loads, but are generally designed for cluster-scale datacenter deployment, without accounting for edge-specific constraints like heterogeneous hardware or limited bandwidth. Complementary to placement strategies, communication-aware MoE systems, such as PipeMoE [shi2023pipemoe], ScheMoE [shi2024schemoe], and TUTEL [hwang2023tutel], aim to improve system efficiency through pipelining, hybrid parallelism, or fine-grained communication scheduling. While primarily designed for training, their techniques offer complementary benefits for distributed inference. In contrast to these prior efforts, DanceMoE targets collaborative inference across memory-limited edge servers. It introduces activation-aware placement and lightweight migration strategies tailored to heterogeneous devices and variable workloads—an underexplored yet increasingly important setting for scalable MoE deployment.